# Explainable AI for computational pathology identifies model limitations and tissue biomarkers

### v1

@@ Abstract
Deep learning models have shown promise in histopathology image analysis, but their opaque decision-making process poses challenges in high-risk medical scenarios. Here we introduce HIPPO, an explainable AI method that interrogates attention-based multiple instance learning (ABMIL) models in computational pathology by generating counterfactual examples through tissue patch modifications in whole slide images. Applying HIPPO to ABMIL models trained to detect breast cancer metastasis reveals that they may overlook small tumors and can be misled by non-tumor tissue, while attention maps—widely used for interpretation—often highlight regions that do not directly influence predictions. By interpreting ABMIL models trained on a prognostic prediction task, HIPPO identified tissue areas with stronger prognostic effects than high-attention regions, which sometimes showed counterintuitive influences on risk scores. These findings demonstrate HIPPO’s capacity for comprehensive model evaluation, bias detection, and quantitative hypothesis testing. HIPPO greatly expands the capabilities of explainable AI tools to assess the trustworthy and reliable development, deployment, and regulation of weakly-supervised models in computational pathology.
@@ Introduction
Digital pathology has emerged as a transformative force in medicine, ushering in an era where computational methods can augment and enhance the diagnostic and prognostic capabilities of pathologists. By digitizing whole slide images (WSIs) of tissue specimens, this field has opened up new avenues for applying advanced machine learning techniques to analyze complex histological patterns and features. The potential impact of computational pathology is far-reaching, promising to improve diagnostic accuracy, standardize interpretation, and uncover novel biomarkers that may inform personalized treatment strategies [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13].
Recently, attention-based multiple instance learning (ABMIL) [14] has emerged as a powerful approach to analyze WSIs for various pathological tasks, demonstrating performance that often rivals or surpasses that of expert pathologists [15]. ABMIL models treat each WSI as a collection of smaller image patches (instances) and use attention mechanisms to identify and focus on the most relevant regions for the task at hand. Importantly, multiple instance learning allows ABMIL models to learn from specimen-level labels, not requiring exhaustive pixel-level annotations, which are time-consuming and costly to obtain[15]. This feature makes ABMIL models particularly well-suited for tasks such as cancer detection [16, 17], diagnosis [18, 19, 20, 21], identification of primary cancer origin [22], grading [17, 23, 24], genomic aberration detection [25, 26, 27, 28], molecular phenotyping [29, 30, 31], treatment response prediction [32, 33, 34], and prognostication [35, 36, 33, 37].
However, the widespread adoption of ABMIL models in clinical settings is hindered by challenges in model interpretability and trustworthiness [38, 39, 9, 10]. A key limitation lies in the heavy reliance of interpretations based on ABMIL’s attention, which is often used as a proxy for understanding model behavior. While attention highlights regions of interest within a WSI, they do not necessarily reflect the direct influence of these regions on model predictions [40, 41]. This disconnect between attention and model output can lead to misinterpretations of model behavior, potentially eroding trust in the model’s decisions and limiting its clinical utility [42, 43, 44, 45]. In addition, post hoc model explanations via attribution methods, such as LIME [46] and SHAP [47], make restrictive additive or linear assumptions of individual pixels, which have been argued to not reflect a model’s decision making process [48].
@@
To address these challenges, we introduce HIPPO (Histopathology Interventions of Patches for Predictive Outcomes), an explainable AI method designed to enhance trust in ABMIL models and provide deeper insights into their decision-making processes. HIPPO goes beyond traditional attention-based interpretations by quantitatively assessing the impact of specific tissue regions on model predictions. By simulating targeted interventions through the occlusion or inclusion of individual or groups of patches, HIPPO enables a more nuanced understanding of how different histological features influence ABMIL model outputs.
We demonstrate the utility of HIPPO by applying it to two major tasks in computational pathology: metastasis detection and cancer prognosis prediction. In the context of metastasis detection, we evaluate five foundation models in pathology using the CAMELYON16 dataset [16]. Our analysis uncovers model-specific limitations and biases that would have remained hidden using attention mechanism alone. We reveal that some models rely heavily on extratumoral tissue for metastasis detection, while others are surprisingly insensitive to small tumor regions. These findings highlight the importance of rigorous model evaluation beyond standard performance metrics and underscore the potential of HIPPO in identifying when and why models might fail.
@@
For cancer prognosis, we apply HIPPO to ABMIL models trained on breast cancer and cutaneous melanoma datasets from The Cancer Genome Atlas. Our results demonstrate that HIPPO can identify tissue regions more strongly associated with prognosis compared to those highlighted by attention. Strikingly, we find that high-attention regions can sometimes have counterintuitive effects on prognostic predictions, further emphasizing the limitations of relying solely on attention for model interpretation. By quantitatively assessing the impact of tumor-infiltrating lymphocytes (TILs) on model predictions, we confirm that the models have captured the known prognostic significance of TILs in both breast cancer and melanoma. This ability to link model behavior to established biological knowledge is crucial for building trust in AI-driven prognostic tools.
@@
Beyond model interpretation, HIPPO opens up new possibilities for in silico experimentation in computational pathology. We showcase this potential by using HIPPO to simulate the effects of autologous TIL therapy in melanoma patients. By digitally replicating TIL-positive regions in specimens with poor prognosis, we demonstrate a proof-of-principle approach for identifying patients who potentially might benefit from this immunotherapy. This application illustrates how HIPPO can bridge the gap between computational predictions and clinically actionable insights.
@@
As computational pathology continues to advance, the need for robust, interpretable, and trustworthy AI models becomes increasingly critical. HIPPO represents a significant step forward in this direction, offering a powerful tool for uncovering the strengths, limitations, and potential biases of ABMIL models in pathology. By providing a more comprehensive understanding of model behavior, HIPPO not only enhances the interpretability of existing models but also paves the way for developing more reliable and clinically relevant AI tools in pathology. As we demonstrate across multiple applications, from metastasis detection to prognostic modeling, HIPPO has the potential to accelerate the translation of computational pathology into clinical practice, ultimately improving patient care and outcomes.
@@ Results
HIPPO is a specimen-level perturbation toolkit that explains weakly-supervised models in computational pathology (Fig. 1a). The fundamental goal of HIPPO is to explore counterfactual (i.e., “what if”) scenarios that are infeasible to realize in actual tissue samples. For instance, it would be impractical to directly manipulate the tumor microenvironment of a tissue specimen to understand its effect on a prognostic model. Instead, we can digitally modify a WSI that simulates this intervention. HIPPO enables in silico interventions through the occlusion or inclusion of single or multiple patches, utilizing the resulting ABMIL model predictions as counterfactual outcomes. HIPPO provides quantitative insights into how specific tissue alterations impact pathological assessments through the lens of the AI model. These assessments can include but are not limited to, patient prognosis, treatment response prediction, metastasis detection, inference of spatial transcriptomics, gene mutation detection, and microsatellite instability identification. Applying HIPPO to ABMIL models enables researchers, regulators, and clinicians to elucidate model behavior and assess the reliability of model outputs in high-risk clinical contexts.
@@
HIPPO-search-low-effect: a greedy search algorithm to identify the regions that lead to the slightest change in model predictions. This can be used to identify the regions that do not change model predictions and, therefore, are not necessary for the output of a model. After removing unnecessary regions, the remaining regions can be considered sufficient for the model prediction (Fig. 1c).
HIPPO generates counterfactual outcomes through the lens of an AI model. Thus, if the AI model has limitations in fully capturing complex biological input-output relationships, HIPPO’s explanations will reflect these limitations. HIPPO provides unprecedented access to interrogating the biological factors underlying prediction, providing insight into an AI’s decision-making process.
@@
In this dataset, expert pathologists finely annotated metastatic regions. This allows us to use HIPPO-knowledge to determine whether metastatic regions are necessary for detecting breast cancer metastasis. Specifically, for patients who were positive for metastasis, we removed the patches that intersected with the tumor annotations, effectively creating a version of the specimen that does not contain metastasis. We compared model predictions before and after the intervention. Specificity was calculated as the ratio of true negatives to all negative samples. In this set of counterfactuals, all specimens were negative, so the specificity represented the proportion of correct negative predictions by the models. Notably, the UNI-based model exhibited the lowest specificity (0.73) in these counterfactual examples despite achieving the highest balanced accuracy on the original test set (1.00). This discrepancy was particularly pronounced in counterfactual specimens that originally contained macrometastases (specificity 0.59), suggesting that the UNI-based ABMIL model uses tissue outside of the tumor region to drive positive metastasis predictions. The REMEDIS-based model exhibited a similar trend, with a specificity of 0.77 in counterfactuals derived from macrometastases. In contrast, the other models showed less dependence on extratumoral tissue (sensitivity of Phikon-based, 0.86; CTransPath-based, 0.92; RetCCL-based, 0.88), indicating that their predictions are primarily driven by tumor epithelial cells rather than other tissue components (Fig. 2c).
We hypothesized that the low specificity of the UNI-based ABMIL model may be attributed to metastasis-induced alterations in the surrounding tumor microenvironment. To investigate this, we used HIPPO-knowledge to remove increasingly larger regions surrounding the tumor annotation progressively and quantified the effect on metastasis detection. As the extent of peritumoral tissue removal increased, the UNI-based model was consistently more likely to predict the absence of metastasis. Specificity increased from 0.73 at dilation of , to 0.78 at , to 0.80 at , to 0.86 at , and to 0.88 at . This was driven primarily by macrometastatic specimens, where specificity increased from 0.59 at dilation of to 0.68 at dilation of , to 0.73 at dilation of , to 0.82 at dilation of , to 0.86 at dilation of . Notably, other ABMIL models remained largely unaffected by peritumoral tissue removal, highlighting a unique characteristic of the UNI-based model (Supplementary Fig. 1). In summary, HIPPO enabled the quantitative exploration of peritumoral tissue on metastasis detection.
@@
While necessity assesses the importance of a feature or feature set, it does not inform whether the feature set is sufficient for model predictions. Metastasis detection models must be able to detect tumor regions no matter how small. Using HIPPO-knowledge, we tested the sufficiency of metastatic regions using two methods: removing all non-tumor patches and measuring model outputs and adding tumor regions to normal specimens and measuring model outputs.
@@
To identify the largest amount of tumor that would go undetected by a model, we also used a HIPPO search algorithm, HIPPO-search-small-effect. We found that in some cases, regions up to could be added into a negative counterfactual while still maintaining a negative detection. Indeed, the tumor patches that were insufficient to drive large effect sizes were largely similar to the sufficient patches, though some insufficient patches contained adipose cells along with tumor epithelial cells (Supplementary Fig. 7). This shows that there exist regions within tumors that would go unseen by an ABMIL-based metastasis detection model. These biases should be explored further prior to the clinical use of metastasis detection models.
Non-tumor tissue can cause false positive metastasis detections
@@
Despite the lack of tumor cells, halos of peritumoral tissue were sufficient to drive a positive metastasis detection. Halos in width beginning from the edge of the expert tumor annotation from 20% (n=10) of positive specimens caused positive predictions in the UNI-based model when embedded in normal specimens, whereas 10% (n=5) caused positive predictions in the Phikon-based model. When starting the halo outside of the tumor annotation, there were 14% (n=7) positive predictions in the UNI-based model and 10% (n=5) for the Phikon-based model. Thinner halos of beginning outside of the tumor annotation also caused positive predictions: 10% (n=5) in the UNI-based model and 6% (n=3) in the Phikon-based model (Supplementary Fig. 8). These results demonstrate that the models learned an association between peritumoral tissue and the presence of metastasis despite the peritumoral tissue not containing any metastasis. This highlights an important bias that could not have been uncovered using attention alone. HIPPO enabled the quantitative assessment of peritumoral tissue on metastasis detection.
@@
Identifying spurious correlations in deep learning models for medical imaging is crucial to ensure reliable and clinically relevant results. To test HIPPO’s ability to identify spurious correlations, we conducted an experiment where we deliberately introduced an artificial bias into the CAMELYON16 dataset (Figs. 4a and 4b). Specifically, blue squares were added to all negative images. This mimics the plausible scenario in which a pathologist marks certain slides with a blue marker. However, in doing so, it introduces a strong spurious correlation with labels. We hypothesized that the models would learn that slides were negative if a blue region was present and that slides lacking this blue region are positive (as blue regions are easier to identify compared to more variable tumor regions).
An ABMIL model was trained on the modified training data using UNI embeddings. The model achieved a balanced accuracy of 1.0 on the test set, suggesting the spurious correlations created a trivial prediction task. By performing standard model interpretation using attention, we found that metastatic regions were considered highly important (Fig. 4c). However, removing these regions using HIPPO did not alter the model predictions, demonstrating that tumor regions were not important for model predictions despite a strong attention assignment. This highlights an important weakness of attention: the disconnect between attended regions and model predictions.
Knowing that the metastatic regions did not affect model outputs, we used the search algorithm HIPPO-search-high-effect to identify the regions that maximally drove positive tumor predictions in both models using one positive specimen, test_002. Given that the model trained with spurious correlations uses the lack of a blue square as a cue for positive specimens, we expected that no individual patches would drive the positive metastasis output and that tumor regions would not have a high effect on the prediction. Indeed, effect sizes were small and evenly distributed across the WSI (minimum , maximum , mean , and median ), indicating that no single region contributed strongly to the model prediction (Fig. 4d). By contrast, applying this search algorithm to the model trained on the original CAMELYON16 dataset, we found that patch effect sizes were higher (minimum , maximum , mean , and median ), and high effect patches were within expert tumor annotations (Fig. 4e). By tying interpretation analysis directly to predictions, HIPPO-based interpretations may provide more reliable explanations of model predictions.
@@
Having demonstrated HIPPO’s effectiveness in metastasis detection, where the regions of interest are well-defined and were previously annotated by expert pathologists, we extended our investigation to the more complex domain of cancer prognosis. Unlike the clear delineation of tumor regions in metastasis detection, prognostic factors in WSIs are multifaceted and less clearly defined. We applied HIPPO to prognostic models that generate risk scores from WSIs, aiming to identify the tissue regions driving these predictions. Our experiments with HIPPO yielded two key insights. First, HIPPO’s search algorithms demonstrated superior ability in identifying tissue patches that consistently and significantly influence risk predictions compared to conventional attention-based methods. While attention mechanisms yielded mixed effects — potentially identifying regions that counterintuitively drive lower risk in otherwise high-risk specimens — HIPPO provided a more consistent, reliable, and quantitative assessment of the regions that drive risk. Second, HIPPO’s unique features enable in silico experiments to measure the effects of targeted tissue interventions on prognostic outcomes through the lens of the ABMIL model. HIPPO’s potential to accelerate the discovery and validation of prognostic tissue biomarkers is an exciting development in cancer research, potentially bridging the gap between computational predictions and clinical actionability.

### v2

@@ Abstract
Introduction: Deep learning models hold great promise for digital pathology, but their opaque decision-making processes undermine trust and hinder clinical adoption. Explainable AI methods are essential to enhance model transparency and reliability.
Methods: We developed HIPPO, an explainable AI framework that systematically modifies tissue regions in whole slide images to generate image counterfactuals, enabling quantitative hypothesis testing, bias detection, and model evaluation beyond traditional performance metrics. HIPPO was applied to a variety of clinically important tasks, including breast metastasis detection in axillary lymph nodes, prognostication in breast cancer and melanoma, and IDH mutation classification in gliomas. In computational experiments, HIPPO was compared against traditional metrics and attention-based approaches to assess its ability to identify key tissue elements driving model predictions.
Results: In metastasis detection, HIPPO uncovered critical model limitations that were undetectable by standard performance metrics or attention-based methods. For prognostic prediction, HIPPO outperformed attention by providing more nuanced insights into tissue elements influencing outcomes. In a proof-of-concept study, HIPPO facilitated hypothesis generation for identifying melanoma patients who may benefit from immunotherapy. In IDH mutation classification, HIPPO more robustly identified the pathology regions responsible for false negatives compared to attention, suggesting its potential to outperform attention in explaining model decisions.
Conclusions: HIPPO expands the explainable AI toolkit for computational pathology by enabling deeper insights into model behavior. This framework supports the trustworthy development, deployment, and regulation of weakly-supervised models in clinical and research settings, promoting their broader adoption in digital pathology.
@@ Introduction
Digital pathology has emerged as a transformative force in medicine, ushering in an era where computational methods can augment and enhance the diagnostic and prognostic capabilities of pathologists. By digitizing whole slide images (WSIs) of tissue specimens, this field has opened new avenues for applying advanced machine learning techniques to analyze complex histological patterns and features. The potential impact of computational pathology is far-reaching, promising to improve diagnostic accuracy, standardize interpretation, and uncover novel biomarkers that may inform personalized treatment strategies [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15].
Building on these advances, the development of pathology foundation models represents a leap forward in computational pathology. These models have significantly expanded the field’s capabilities by providing powerful, generalizable tools for analyzing histological features [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. They are trained on vast collections of histology image patches in a self-supervised manner, enabling them to learn general-purpose representations of histological features. By focusing on patches, foundation models can efficiently process the large size and complexity of WSIs while capturing key local patterns. The representations learned by these models can be utilized in downstream tasks by incorporating patch-level embeddings as inputs for downstream models. For specimen-level prediction tasks, these embeddings are often integrated into an attention-based multiple instance learning (ABMIL[28]) framework. In this approach, the foundation model provides patch-level feature representations, and ABMIL aggregates and prioritizes relevant patches to generate specimen-level predictions. This combination has quickly become the standard for leveraging foundation models in specimen-level classification, offering a scalable and effective way to utilize their pre-trained knowledge for complex clinical and research applications.
Unlike traditional image classification methods, ABMIL models rely on specimen-level labels, eliminating the need for exhaustive pixel-level annotations [29]. This flexibility has enabled ABMIL to excel across a range of tasks, including cancer detection [30, 31], diagnosis [32, 33, 34, 35], identification of primary cancer origin [36], grading [31, 37, 38], genomic aberration detection [39, 40, 41, 42], molecular phenotyping [43, 44, 45], treatment response prediction [46, 47, 48], and prognostication [49, 50, 47, 51]. The integration of foundation models with ABMIL has proven particularly effective, combining the generalization capabilities of pre-trained models with the task-specific focus of ABMIL to advance computational pathology.
However, the widespread adoption of ABMIL models in clinical settings is hindered by challenges in model interpretability and trustworthiness [52, 53, 9, 10, 54, 55, 56]. Explainable AI methods have been proposed to address these challenges by providing insights into model behavior and identifying features that may influence predictions. Techniques such as LIME [57], SHAP [58], and SmoothGrad [59] offer strategies to estimate feature contributions in input images. However, these methods often rely on linear or additive approximations, which can oversimplify complex model behaviors and fail to align with actual model predictions [60, 61].
Attention mechanisms have emerged as a widely used approach for interpreting ABMIL models, as evidenced by their extensive use in recent studies [28, 46, 36, 50, 62, 63, 22, 64, 45, 65, 66, 67, 68, 69]. Although attention highlights regions of interest within WSIs, it does not directly quantify the influence of these regions on model predictions [70, 71]. This gap between attention and predictive influence can lead to misinterpretation, undermining trust in model decisions and limiting clinical applicability [72, 73, 74, 75]. Addressing this limitation requires methods that explicitly evaluate the contribution of specific regions to model outputs.
One such approach was introduced by Zeiler and Fergus, occluding parts of an image with gray boxes and observing how the model’s predictions change [76]. This approach can help identify which parts of an image are most relevant to a model’s decision. However, when applied to pathology, this method has limitations. Covering regions of WSIs with gray patches can cause erratic behavior in the model, as these unnatural patterns are not encountered during training, leading to interpretations that may be difficult to trust. While identifying regions to occlude in natural images often aligns with human intuition, this becomes more challenging in WSIs, where multiple regions of interest may collectively influence predictions. Additionally, occlusion methods only indicate the importance of regions but do not address whether the identified features are sufficient to drive the model’s predictions.
To better understand how specific tissue features influence model predictions, generating realistic counterfactual WSIs is essential. Counterfactual WSIs allow us to explore “what if” scenarios by systematically altering tissue features and observing the impact on model behavior. For example, resizing a tumor region could help determine the model’s sensitivity to tumor size, while modifying the number and distribution of tumor-infiltrating lymphocytes might reveal their role in influencing predictions, such as for prognosis or treatment response.
However, creating reliable counterfactual WSIs presents significant challenges. First, precise segmentation of tissue structures is necessary to enable targeted alterations without affecting unrelated regions. Second, robust inpainting techniques are required to fill in altered regions realistically, ensuring the resulting WSI appears biologically plausible. Although recent advances in generative AI show promise for creating synthetic images [77], applying these methods at the gigapixel scale for WSIs remains a technical challenge. The ability to generate realistic and valid counterfactual WSIs for clinical use has yet to be proven.
@@
To address these challenges, we introduce HIPPO (Histopathology Interventions of Patches for Predictive Outcomes), an explainable AI method designed to enhance trust in ABMIL models and provide deeper insights into their decision-making processes. HIPPO goes beyond traditional attention-based interpretations by quantitatively assessing the impact of specific tissue regions on model predictions. HIPPO also leverages properties of multiple instance learning models to generate valid WSIs through the occlusion or inclusion of individual or groups of patches, simulating targeted interventions to understand how different histological features influence ABMIL model outputs.
@@
We validate HIPPO across three distinct tasks: metastasis detection, prognostication, and IDH mutation classification. In the context of metastasis detection, we evaluate five foundation models in pathology using the CAMELYON16 dataset [30]. This dataset includes expert-annotated labels and thus allows us to rigorously assess HIPPO’s ability to align with established ground truth and demonstrate its practical utility. Our analysis uncovers model-specific limitations and biases that would have remained hidden using attention mechanism alone. We reveal that some models rely heavily on extratumoral tissue for metastasis detection, while others are surprisingly insensitive to small tumor regions. These findings highlight the importance of rigorous model evaluation beyond standard performance metrics and underscore the potential of HIPPO in identifying when and why models might fail. Building on this validation, we then aplied HIPPO to cancer prognostication in breast cancer and cutaneous melanoma datasets from The Cancer Genome Atlas. Our results demonstrate that HIPPO can identify tissue regions more strongly associated with prognosis compared to those highlighted by attention. Strikingly, we find that high-attention regions can sometimes have counterintuitive effects on prognostic predictions, further emphasizing the limitations of relying solely on attention for model interpretation. By quantitatively assessing the impact of tumor-infiltrating lymphocytes (TILs) on model predictions, we confirm that the models have captured the known prognostic significance of TILs in both breast cancer and melanoma. This ability to link model behavior to established biological knowledge is crucial for building trust in AI-driven prognostic tools. Finally, we applied HIPPO to IDH mutation classification in glioma and found that HIPPO outperformed attention in identifying the histological regions driving false negatives. By providing more precise and interpretable insights into model predictions, HIPPO enabled a clearer understanding of the reasons behind misclassifications.
@@
Beyond model interpretation, HIPPO opens new possibilities for virtual hypothesis generation in computational pathology. We showcase this potential by using HIPPO to simulate the effects of autologous TIL therapy in melanoma patients. By digitally replicating TILs in specimens with poor prognosis, we demonstrate a proof-of-principle approach for identifying patients who potentially might benefit from this immunotherapy. This application illustrates how HIPPO can bridge the gap between computational predictions and clinically actionable insights.
@@
As computational pathology continues to advance, the need for robust, interpretable, and trustworthy AI models becomes increasingly critical. HIPPO represents a significant step forward in this direction, offering a powerful tool for uncovering the strengths, limitations, and potential biases of ABMIL models in pathology. By providing a more comprehensive understanding of model behavior, HIPPO not only enhances the interpretability of existing models but also paves the way for developing more reliable and clinically relevant AI tools in pathology. As we demonstrate across multiple applications, from metastasis detection to prognostic modeling and mutation classification, HIPPO has the potential to accelerate the translation of computational pathology into clinical practice, ultimately improving patient care and outcomes.
@@ Results
HIPPO is a specimen-level perturbation toolkit that explains weakly-supervised models in computational pathology (Fig. 1a). The fundamental goal of HIPPO is to explore counterfactual (i.e., “what if”) scenarios that are infeasible to realize in actual tissue samples. For instance, it would be impractical to directly manipulate the tumor microenvironment of a tissue specimen to understand its effect on a prognostic model. Instead, we can digitally modify a WSI that simulates this intervention. HIPPO enables virtual interventions through the occlusion or inclusion of single or multiple patches, utilizing the resulting ABMIL model predictions as counterfactual outcomes. HIPPO provides quantitative insights into how specific tissue alterations might impact pathological assessments through the lens of the AI model. These assessments can include but are not limited to, patient prognosis, treatment response prediction, metastasis detection, inference of spatial transcriptomics, gene mutation detection, and microsatellite instability identification. Applying HIPPO to ABMIL models enables researchers, regulators, and clinicians to elucidate model behavior and to better understand the relationship between patch-level tissue characteristics and model outputs.
@@
HIPPO-search-low-effect: a greedy search algorithm to identify the regions that lead to the smallest change in model predictions. This can be used to identify the regions that do not change model predictions and, therefore, are not necessary for the output of a model. After removing unnecessary regions, the remaining regions can be considered sufficient for the model prediction (Fig. 1c).
Hypothesis-driven interventions leverage prior knowledge of relevant patches, which can be identified with the assistance of pathologists or annotation tools. In cases where prior information is unavailable or when important features are sub-visual (e.g., in prognostic prediction tasks), data-driven interventions can automatically identify important patches de novo. Unlike traditional attribution methods that rely on additive approximations, HIPPO supports multi-patch interventions, enabling the exploration of how sets of patches collectively drive non-linear effects on predictions.
It is important to note that HIPPO generates counterfactual predictions based on the capabilities of the underlying AI model. Consequently, if the AI model has limitations in capturing complex biological relationships, these limitations will inherently influence HIPPO’s explanations. Furthermore, the resolution of HIPPO’s interventions is constrained to patches, which may restrict its ability to capture finer-grained details. Despite these constraints, HIPPO offers a novel and powerful approach for interrogating the biological factors driving predictions, shedding light on the decision-making process of AI models in unprecedented ways.
@@
In this dataset, expert pathologists have produced high resolution annotations of metastatic tumor regions. This allows us to use the HIPPO-knowledge method to determine whether metastatic regions are necessary for detecting breast cancer metastasis. Specifically, for patients who were positive for metastasis, we created counterfactual examples by removing all patches that intersected with the tumor annotations. If we assume that a model must rely on tumor-containing patches to make a prediction, an accurate model would always return a negative prediction when all tumor containing patches have been removed. In the following, we use specificity to quantify model behavior. Specificity is defined as the ratio of true negatives to all negative samples. In this set of counterfactuals, all specimens were negative, so the specificity represented the proportion of correct negative predictions by the models.
We compared model predictions before and after removing patches containing tumor. Notably, the UNI-based model exhibited the lowest specificity (0.73) in these counterfactual examples despite achieving the highest balanced accuracy on the original test set (1.00). This discrepancy was particularly pronounced in counterfactual specimens that originally contained macrometastases (specificity 0.59), suggesting that the UNI-based ABMIL model uses tissue outside of the tumor region to drive positive metastasis predictions. The REMEDIS-based model exhibited a similar trend, with a specificity of 0.77 in counterfactuals derived from macrometastases. In contrast, the other models showed less dependence on extratumoral tissue (sensitivity of Phikon-based, 0.86; CTransPath-based, 0.92; RetCCL-based, 0.88), indicating that their predictions are primarily driven by tumor epithelial cells rather than other tissue components (Fig. 2c).
We hypothesized that the low specificity of the UNI-based ABMIL model may be attributed to metastasis-induced alterations in the surrounding tumor microenvironment. To investigate this, we used HIPPO-knowledge to remove increasingly larger regions surrounding the tumor annotation and quantified the effect on metastasis detection. As the extent of peritumoral tissue removal increased, the UNI-based model was consistently more likely to predict the absence of metastasis. Specificity increased from 0.73 at dilation of , to 0.78 at , to 0.80 at , to 0.86 at , and to 0.88 at . This was driven primarily by macrometastatic specimens, where specificity increased from 0.59 at dilation of to 0.68 at dilation of , to 0.73 at dilation of , to 0.82 at dilation of , to 0.86 at dilation of . Notably, other ABMIL models remained largely unaffected by peritumoral tissue removal, highlighting a unique characteristic of the UNI-based model (Supplementary Fig. 1). In summary, HIPPO enabled the quantitative exploration of peritumoral tissue on metastasis detection.
@@
While necessity assesses the importance of a feature or feature set, it does not inform whether the feature set is sufficient for model predictions. Metastasis detection models must be able to detect tumor regions no matter how small. Using the HIPPO-knowledge method, we tested the sufficiency of metastatic regions using two methods: removing all non-tumor patches and measuring model outputs and adding tumor regions to normal specimens and measuring model outputs.
@@
To identify the largest amount of tumor that would go undetected by a model, we also used a HIPPO search method, HIPPO-search-low-effect. We found that in some cases, regions up to could be added into a negative counterfactual while still maintaining a negative detection. Indeed, the tumor patches that were insufficient to drive large effect sizes were largely similar to the sufficient patches, though some insufficient patches contained adipose cells along with tumor epithelial cells (Supplementary Fig. 7). This shows that there exist regions within tumors that would go unseen by an ABMIL-based metastasis detection model. These biases should be explored further prior to the clinical use of metastasis detection models.
Peritumoral tissue can cause false positive metastasis detections
@@
Halos of peritumoral tissue were sufficient to drive positive metastasis detection. Halos in width beginning from the edge of the expert tumor annotation from 20% (n=10) of positive specimens caused positive predictions in the UNI-based model when embedded in normal specimens, whereas 10% (n=5) caused positive predictions in the Phikon-based model. When starting the halo outside of the tumor annotation, there were 14% (n=7) positive predictions in the UNI-based model and 10% (n=5) for the Phikon-based model. Thinner halos of beginning outside of the tumor annotation also caused positive predictions: 10% (n=5) in the UNI-based model and 6% (n=3) in the Phikon-based model (Supplementary Fig. 8). These results demonstrate that the models learned an association between peritumoral tissue and the presence of metastasis despite the peritumoral tissue not containing any metastasis. This highlights an important bias that could not have been uncovered using attention alone. HIPPO enabled the quantitative assessment of peritumoral tissue on metastasis detection in multiple foundation models.
@@
Identifying spurious correlations in deep learning models for medical imaging is crucial to ensure reliable and clinically relevant results. To test HIPPO’s ability to identify spurious correlations, we conducted an experiment where we deliberately introduced an artificial bias into the CAMELYON16 dataset (Supplementary Figs. 9a and 9b). Specifically, blue squares were added to all negative images. This mimics the plausible scenario in which a pathologist marks certain slides with a blue marker. However, in doing so, it introduces a strong spurious correlation with labels. We hypothesized that the models would learn that slides were negative if a blue region was present and that slides lacking this blue region are positive (as blue regions are easier to identify compared to more variable tumor regions).
An ABMIL model was trained on the modified training data using UNI embeddings. The model achieved a balanced accuracy of 1.0 on the test set, suggesting the spurious correlations created a trivial prediction task. By performing standard model interpretation using attention, we found that metastatic regions were considered highly important (Supplementary Fig. 9c). However, removing these regions using HIPPO did not alter the model predictions, demonstrating that tumor regions were not important for model predictions despite a strong attention assignment. This highlights an important weakness of attention: the disconnect between attended regions and model predictions.
Knowing that the metastatic regions did not affect model outputs, we used the search algorithm HIPPO-search-high-effect to identify the regions that maximally drove positive tumor predictions in both models using one positive specimen, test_002. Given that the model trained with spurious correlations uses the lack of a blue square as a cue for positive specimens, we expected that no individual patches would drive the positive metastasis output and that tumor regions would not have a high effect on the prediction. Indeed, effect sizes were small and evenly distributed across the WSI (minimum , maximum , mean , and median ), indicating that no single region contributed strongly to the model prediction (Supplementary Fig. 9d). By contrast, applying this search algorithm to the model trained on the original CAMELYON16 dataset, we found that patch effect sizes were higher (minimum , maximum , mean , and median ), and high effect patches were within expert tumor annotations (Supplementary Fig. 9e). By tying interpretation analysis directly to predictions, HIPPO-based interpretations may provide more reliable explanations of model predictions.
@@
Having demonstrated HIPPO’s effectiveness in metastasis detection, where the regions of interest are well-defined and were previously annotated by expert pathologists, we extended our investigation to the more complex domain of cancer prognosis. Unlike the clear delineation of tumor regions in metastasis detection, prognostic factors in WSIs are multifaceted and less clearly defined. We applied HIPPO to prognostic models that generate risk scores from WSIs, aiming to identify the tissue regions driving these predictions. Our experiments with HIPPO yielded two key insights. First, HIPPO’s search algorithms demonstrated superior ability in identifying tissue patches that consistently and significantly influence risk predictions compared to conventional attention-based methods. While attention mechanisms yielded mixed effects — potentially identifying regions that counterintuitively drive lower risk in otherwise high-risk specimens — HIPPO provided a more consistent, reliable, and quantitative assessment of the regions that drive risk. Second, HIPPO’s unique features enable virtual experiments to measure the effects of targeted tissue interventions on prognostic outcomes through the lens of the ABMIL model. HIPPO’s potential to accelerate the discovery and validation of prognostic tissue biomarkers is an exciting development in cancer research, potentially bridging the gap between computational predictions and clinical actionability.